OpenStudy (anonymous):

6 N acts on an object at 0 deg and 8 N acts at 90 deg. What is the size and direction of their equilibriant force?

OpenStudy (anonymous):

Pythagorean's Theorem is your friend. \[|F| = \sqrt{(F_x ^2 + F_y^2)}\]Realize that because the forces acts at 90 degrees to each other, they represent x and y components. F_x = 6 F_y = -8

OpenStudy (anonymous):

I forgot to include how to calculate direction. This comes from trig. \[\theta = \arctan \left(y \over x \right)\]